two numbers are in the ration 3:8. If the first number is increased by 4 and the second number is decreased by 6, the resulting numbers are in the ratio of 1:2. Find the original numbers.
Answer by checkley77(12844)

(3x+4):(8x-6)=1:2
2(3x+4)=(8x-6)
6x+8=8x-6
-2x=-14
x=-14/-2
x=7
Proof:
(3*7+4):(8*7-6)=1:2
(21+4):(56-6)=1:2
25:50=1:2
1:2=1:2